Ask
your own question, for FREE!
Mathematics
11 Online
A fair coin is tossed twice. what is the probability of obtaining at least one head?
Still Need Help?
Join the QuestionCove community and study together with friends!
There are 4 possibilities HH HT TH TT 3 of these have at least one H ->P = 3/4
You could also recognize, P(at least 1 head) = 1 - P(no heads) = 1 - P( 2 tails) = 1 - 1/2 x 1/2 = 1-1/4 = 3/4
What lokisan is doing is finding the probability that you will get tails twice, (which would be multiplying one half by one half), then subtracting that from one, which would get the probability that you wouldn't get tails twice, which is the same thing as getting heads at least once.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
4 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal